# Sailor: Automating Distributed Training over Dynamic, Heterogeneous, and Geo-distributed Clusters

### v1

@@ 5.2.3. Geo-distributed setups
In Figure 10, we evaluate Sailor’s throughput in a setup with 2 cloud regions and 5 zones (4 zones in us-central1 and 1 zone in us-west1), considering A100 GPUs for the OPT-350M model. We compare Sailor with DTFM, with the exhaustive search to automatically discover parallelization plans. We report training throughput and monetary cost per iteration, taking both the computation and communication cost into account.
DTFM cannot fully leverage multiple zones, mainly due to its suboptimal cost function, and lack of memory footprint estimation. DTFM ranks solutions based on the time spent in data and pipeline parallel communication, which leads to suboptimal solution ranking. Sailor achieves 5.9$\times$ higher throughput and 9.48$\times$ lower cost per iteration. Sailor employs larger microbatch sizes and tensor parallelism degrees, reducing the pipeline and data parallel data transfers. Furthermore, Sailor finds configurations within 1 second, compared to DTFM that needs hundrend of seconds with large clusters (due to the exhaustive search). On GPT-Neo, DTFM fails due to OOMs, while Sailor finds valid plans with a throughput of 0.07–0.21 iters/sec across cluster sizes.
Finally, comparing Sailor’s throughput for the OPT-350M model in the heterogeneous setups with more V100 (Figure 8(b)) and the geo-distributed A100-only setup (Figure 10), shows that the geo-distributed setup achieves up to 2$\times$ higher throughput, and also lower cost.
@@
Figure 10. Throughput of geo-distributed planners with A100-40GB for the OPT-350M model in 5 zones (2 regions)
@@ 5.2.4. Optimization with constraints
Scenario 2: Maximizing throughput with cost constraint of 1.2 USD/iteration: Figure 12 shows the throughput (bars) and iteration cost (asterisks) for this scenario. Most baselines will use all available resources (e.g. all 128 A100 and 128 V100), even if they do not benefit throughput. DTFM does not find a solution as it outputs plans with low throughput and high costs. Sailor outperforms all baselines leading to 1.65-3$\times$ higher throughput while remaining within the cost constraint. Sailor’s plan includes 256 A100 GPUs in two zones, with tensor parallelism of 4, and data parallelism of 64.
@@ 5.3. Sailor scalability study
We evaluated Sailor’s search time varying the number of GPUs and zones or regions with a homogeneous GPU type. The search time remains below 1.5 sec even with 5 GCP zones and 256 A100/zone for the GPT-Neo-2.7B model. In contrast, adding more GPU types has a much higher impact in Sailor’s search time: considering 256 GPUs/type in a single zone, Sailor’s search time is 0.3, 6.2, 4900 sec for 1, 2, 3 GPU types, respectively. Nevertheless, Sailor’s search process is much more efficient than the rest of the heterogeneous baselines: Metis (Um24Metis, ) needs hours to complete its search even with 2 GPU types, while FlashFlex (yan2024flashflex, ) cannot find a valid configuration for any of these setups.
@@ 5.4. Sailor Planner optimization breakdown
Table 3 shows Sailor’s planner search time breakdown when optimizing for throughput, and budget constraint overhead, considering 128 A100 and 128 V100 GPUs for the GPT-Neo-2.7b model. The dynamic-programming-only approach needs hours to complete the search, even in the single-GPU-type case. Introducing the heuristics H1-H3 (4.2.1), which apply in this scenario, dramatically decreases the search time to a couple of seconds. The additional cost constraint increases the search time (4$\times$ increase in the 2-GPU-type scenario) due to the extra iterations caused by the straggler approximations, as described in 4.2.3.
@@ 6. Other Related Work
Asynchronous geo-distributed training: Several systems propose training over geo-distributed, preemptible, and heterogeneous resources by introducing asynchrony and reducing communication via quantization and sparsification. DiLoCo (douillard2024diloco, ) uses federated averaging to reduce communication, performing more local computation before synchronization. SWARM (ryabinin2023swarm, ) proposes a decentralized, model-parallel approach to deal with poorly connected and unreliable devices. CocktailSGD (Wang23cocktailsgd, ) uses gradient compression to improve communication over low-bandwidth networks. These approaches influence training dynamics and are orthogonal to our work. Sailor employs synchronous training, which is preferred in large-scale training (wagenlander24tenplex, ).
Automatic VM selection: CherryPick (Alipourfard17Cherrypick, ), RAMBO (Li21Rambo, ), and PARIS (Yadwadkar17Paris, ) apply Bayesian Optimization or performance modeling to recommend optimal VM types. SkyPilot (Zongheng23Skypilot, ) picks cost-efficient VMs across providers based on workload and user constraints. These methods treat workloads as black boxes and are not tailored for ML training. Srifty  (Luo22Srifty, ), Cynthia (Zheng19Cynthia, ), SpotDNN (Shang2023spotDNN, ), DeepSpot (Lee17DeepSpotCloud, ) find optimal configurations of on-demand and spot instances for ML jobs. However, they target small data-parallel jobs, and are inadequate for large-scale training that involves various types of communication, which is our focus.
@@ 7. Conclusion
We propose Sailor, a system for efficient large-scale training over heterogeneous resources with dynamic availability. Sailor co-optimizes the resource allocation and parallelization plan for a training job to optimize a user-defined objective, under constraints. By combining accurate iteration time and memory estimation, dynamic-programming based search, and domain-specific heuristics, Sailor efficiently navigates the large search space of possible job configurations. Sailor’s distributed training framework supports heterogeneous setups and provides seamless elasticity. Sailor achieves 1.1-5.9$\times$ higher throughput than baselines across homogeneous, heterogeneous, and geo-distributed settings.

### v2

@@ 5.2.3. Geo-distributed setups
In Figures 11 and 12, we evaluate Sailor’s throughput in geo-distributed setups, considering A100 GPUs for the OPT-350M model. We compare Sailor with DTFM, with the exhaustive search to automatically discover parallelization plans. We report training throughput and monetary cost per iteration, taking both the computation and communication cost into account.
Small-scale results on real hardware: Figure 11 shows an experiment with small-scale cluster in 4 cloud zones (2 regions) using 4 and 8 A100 GPUs per zone. DTFM cannot fully leverage multiple zones, mainly due to its suboptimal cost function and lack of memory footprint estimation. DTFM ranks solutions based on the time spent in data and pipeline parallel communication, which leads to suboptimal solution ranking. Furthermore, it uses all cloud regions, which increases communication bottlenecks and cost without increasing throughput. In contrast, Sailor uses only 1 region with all available zones (us-central1), as incorporating an additional region (us-west1) does not improve throughput. Sailor leads to 1.9 and 2.45 higher throughput than DTFM for two cluster sizes. Our simulator was within 3.7% of the real throughput in this scenario.
Figure 11. Throughput of geo-distributed planners with A100-40GB for the OPT-350M model in 4 zones (2 regions) using real GPUs.
Figure 12. Throughput of geo-distributed planners with A100-40GB for the OPT-350M model in 5 zones (2 regions) using our simulator.
Large-scale results on simulation: In Figure 12 we use our simulator to evaluate larger cluster sizes with 5 cloud zones (2 regions). Sailor achieves 5.9 higher throughput and 9.48 lower cost per iteration than DTFM. Sailor employs larger microbatch sizes and tensor parallelism degrees, reducing the pipeline and data parallel data transfers. Furthermore, Sailor finds configurations within 1 second, compared to DTFM that needs hundrend of seconds with large clusters (due to the exhaustive search). On GPT-Neo, DTFM fails due to OOMs, while Sailor finds valid plans with a throughput of 0.07–0.21 iters/sec across cluster sizes.
Finally, comparing Sailor’s throughput for the OPT-350M model in the heterogeneous setups with more V100 (Figure 8(b)) and the geo-distributed A100-only setup (Figure 12), shows that the geo-distributed setup achieves up to 2 higher throughput, and also lower cost.
@@ 5.2.4. Optimization with constraints
Scenario 2: Maximizing throughput with cost constraint of 1.2 USD/iteration: Figure 14 shows the throughput (bars) and iteration cost (asterisks). Most baselines will use all available resources (e.g. all 128 A100 and 128 V100), even if they do not benefit throughput. DTFM does not find a solution as it outputs plans with low throughput and high costs. Sailor outperforms all baselines leading to 1.65-3 higher throughput while remaining within the cost constraint. Sailor’s plan includes 256 A100 GPUs in two zones, with tensor parallelism of 4, and data parallelism of 64.
@@ 5.3. Sailor scalability study
We evaluated Sailor’s search time varying the number of GPUs and zones or regions with a homogeneous GPU type. The search time remains below 1.5 sec even with 5 GCP zones and 256 A100/zone for the GPT-Neo-2.7B model. In contrast, adding more GPU types has a much higher impact in Sailor’s search time: considering 256 GPUs/type in a single zone, Sailor’s search time is 0.3, 6.2, 4900 sec for 1, 2, 3 GPU types, respectively. Nevertheless, Sailor’s search process is much more efficient than the rest of the heterogeneous baselines: Metis (Um et al., 2024) needs hours to complete its search even with 2 GPU types, while FlashFlex (Yan et al., 2024) cannot find a valid configuration for any of these setups.
@@ 5.4. Sailor Planner optimization breakdown
Table 3 shows Sailor’s planner search time breakdown when optimizing for throughput, and budget constraint overhead, considering 128 A100 and 128 V100 GPUs for the GPT-Neo-2.7b model. The dynamic-programming-only approach needs hours to complete the search, even in the single-GPU-type case. Introducing the heuristics H1-H3 (4.2.1), which apply in this scenario, dramatically decreases the search time to a couple of seconds. The additional cost constraint increases the search time (4 increase in the 2-GPU-type scenario) due to the extra iterations caused by the straggler approximations, as described in 4.2.3.
@@
5.5. Reconfiguration overheads
We measure Sailor’s reconfiguration time on a cluster of 16 V100 GPUs for the OPT-350M model. When 4 more GPUs are added, the controller re-invokes the planner, instructs workers to clean up (e.g., destroy NCCL groups), and broadcasts the new plan and topology. Planning takes 0.1 seconds, process cleanup takes 3 seconds, and topology broadcast (using grpc) takes 1.25 seconds. After the workers have received the new plan and topology, they reinitialize NCCL communication groups (for data, pipeline, tensor parallelism) in 4.5 seconds, redefine the model and optimizer in 2 seconds, redefine the dataloaders in 0.5 seconds, and resume training. While moderate at this scale, NCCL initialization can take minutes on thousands of GPUs (Jiang et al., 2024a). Existing methods to reduce this overhead (Jiang et al., 2024a) can be integrated into Sailor.
6. Discussion
Planner and simulator limitations: Our planner and simulator currently support only the 1F1B pipeline parallel schedule, and do not incorporate optimizations such as activation offloading (Ren et al., 2021) or rematerialization (Yuan et al., 2024). Adding support for these optimizations is left for future work.
Additional challenges with heterogeneous hardware: Training over heterogeneous and geo-distributed datacenters can introduce additional challenges. Heterogeneity in accelerator vendors (e.g., NVIDIA vs. AMD) and network links (e.g., Infiniband vs. Ethernet) may prevent the use of high-performance collective communication libraries (e.g., NCCL), which often assume uniform network protocols. To maximize performance, collectives must be adapted to heterogeneous links. Furthermore, geo-distributed networks are prone to unreliability, including unpredictable jitter and packet loss. Training frameworks and algorithms should detect and adapt to such issues. Finally, even though Sailor optimizes parallelization strategies, it is based on strategies that were first introduced for homogeneous settings (e.g. 1F1B pipeline schedule). Achieving high performance in such contexts may require developing new schedules that more effectively overlap computation and communication, thereby reducing bubble times (Chen et al., 2025).
@@ 7. Other Related Work
Asynchronous geo-distributed training: Several systems propose training over geo-distributed, preemptible, and heterogeneous resources by introducing asynchrony and reducing communication via quantization and sparsification. DiLoCo (Douillard et al., 2024) uses federated averaging to reduce communication, performing more local computation before synchronization. SWARM (Ryabinin et al., 2023) proposes a decentralized, model-parallel approach to deal with poorly connected and unreliable devices. CocktailSGD (Wang et al., 2023) uses gradient compression to improve communication over low-bandwidth networks. These approaches influence training dynamics and are orthogonal to our work. Sailor employs synchronous training, which is preferred in large-scale training (Wagenländer et al., 2024).
Automatic VM selection: CherryPick (Alipourfard et al., 2017), RAMBO (Li et al., 2021), and PARIS (Yadwadkar et al., 2017) apply Bayesian Optimization or performance modeling to recommend optimal VM types. SkyPilot (Yang et al., 2023) picks cost-efficient VMs across providers based on workload and user constraints. These methods treat workloads as black boxes and are not tailored for ML training. Srifty  (Luo et al., 2022), Cynthia (Zheng et al., 2019), SpotDNN (Shang et al., 2023), DeepSpot (Lee and Son, 2017) find optimal configurations of on-demand and spot instances for ML jobs. However, they target small data-parallel jobs, and are inadequate for large-scale training that involves various types of communication, which is our focus.
Heterogeneous and Geo-distributed ML Inference: Recent works have proposed systems for ML inference on heterogeneous and geo-distributed resources (Mei et al., 2025; Jiang et al., 2024b; Mao et al., 2025). Although training and inference workloads differ significantly (e.g. in runtime and memory footprint estimation), synchronization and communication patterns as well as planning decisions such as partitioning models across resources are relevant to both. Helix (Mei et al., 2025) and HexGen (Jiang et al., 2024b) consider both heterogeneous and geo-distributed resources for LLM inference. Helix uses a time-consuming MILP algorithm for model placement, which is not suitable for environments with high resource availability. HexGen uses a more lightweight dynamic programming approach for splitting tensor and pipeline parallelism across resources, making it more appropriate for dynamic settings. However, both of these works optimize only for performance, ignoring the monetary communication costs that arise in geo-distributed settings. SkyServe (Mao et al., 2025) places replicas of models across zones and regions, but restricts each replica to a single zone and homogeneous resources. As replicas do not communicate in inference, inter-zone communication is not considered. In contrast, Sailor targets geo-distributed training, which requires frequent communication among data-parallel replicas, substantially increasing scheduling complexity.
@@ 8. Conclusion
We propose Sailor, a system for efficient large-scale training over heterogeneous resources with dynamic availability. Sailor co-optimizes the resource allocation and parallelization plan for a training job to optimize a user-defined objective, under constraints. By combining accurate iteration time and memory estimation, dynamic-programming based search, and domain-specific heuristics, Sailor efficiently navigates the large search space of possible job configurations. Sailor’s distributed training framework supports heterogeneous setups and provides seamless elasticity. Sailor achieves 1.1-5.9 higher throughput than baselines across homogeneous, heterogeneous, and geo-distributed settings.
9. Acknowledgements
We thank the SOSP’25 anonymous reviewers and our shepherd, Ionel Gog, for their insightful feedback. We also thank Christina Giannoula for her feedback on the paper, and Michal Friedman for the helpful discussions. This work was supported under project ID infra02 as part of the Swiss AI Initiative, through a grant from the ETH Domain and computational resources provided by the Swiss National Supercomputing Centre (CSCS) under the Alps infrastructure. We thank the CSCS team for their technical support. Foteini Strati is supported by the Swiss National Science Foundation (project number 200021_204620). George Manos is an Onassis Foundation scholar.